a body of mass m, moving along the positive x direction is subjected to a resistive force F = -ku^2, where k is a constant and u is the particle velocity. If m = 10 kg, u = 10 m/s at t = 0s and k = 2N the velocity when t = 2s is

Grade:

1 Answers

Eshan
5 years ago
Dear student,

F=ma=m\dfrac{du}{dt}=-ku^2
\implies (\dfrac{1}{u}-\dfrac{1}{u_0})=\dfrac{k}{m}t
\implies (\dfrac{1}{u}-\dfrac{1}{10})=\dfrac{2}{10}(2)
\implies u=2m/s